OpenStudy (anonymous):

Number Relation Problem. The sum of the digits of a two-digit number is 13.If the digits are reversed, the resulting number is 9 less than the original number. What is the original number?

OpenStudy (anonymous):

x+y=13 10x+y=10y+x+9 9x=9y+9 x=y+1 2y+1 =13 y=6 x+7

OpenStudy (anonymous):

x=7 that is

OpenStudy (anonymous):

this is a final answers

OpenStudy (anonymous):

The number is \(76\).

OpenStudy (anonymous):

Consider the number to be of the form XY, where x and y are both numbers between 0 and 9. The number when reversed is 9 less than the original number implies that \(x>y\). So, we will be left with the following choices for the number: \(94, 85\) or \(76\). Find the difference between each number and its reverse, and you will find that that \(76-67=9\).

OpenStudy (anonymous):

cant you solve the problem

OpenStudy (anonymous):

I just did!! The original number is \(76\).

OpenStudy (anonymous):

It's a two-digit number, the sum of the digits \(7+6\) is \(13\) and .If the digits are reversed, the resulting number \(67\) is 9 less than the original number \(76\).

OpenStudy (anonymous):

I believe you made a mistake somewhere lalaly :) Let the original number be \(10x+y=10x+13-x=9x+13\) then the reversed number is \(10y+x=10(13-x)+x=130-9x\) And since the original number is 9 greater than the reversed one, we have this equation \(9x+13=130-9x+9 \implies x=7\). But \(y=13-7=6\). And therefore the original number is \(76\). Did I make a mistake somewhere?!